Write the equation of a line that is perpendicular to the given line and that passes through the given point.
y= 23x+9;(–6,5)

a. y=−23x+1 c. y=23x+9
b. y=−3x+3 d. y=−3x−4 222

Answer by mananth(16946)

y= 23x+9;(–6,5)

y= 23x+9;(–6,5)

slope of line = 23
slope of line perpendicular to this line = -1/23
passes through (-6,5)
the general equation
(y-y1)= m(x-x1)
(y-5)=(-1/23)(x-(-6))
multiply equation by 23
23(y-5)=-1(x+6)
23y-115=-x-6
23y+x=115+6
x+23y=121
write in the form required